Please help me figure out this problem. The perimeter of a square porch is 3 more than the area. Find the length of a side, it's area, and it's perimeter.
Found 2 solutions by stanbon, Theo:
Answer by stanbon(75887)

The perimeter of a square porch is 3 more than the area. Find the length of a side, it's area, and it's perimeter.
-------------
Perimeter = 4s where s is an edge of the square
Area = s^2
-----
Equation:
4s = s^2 + 3
-------------------
s^2 - 4s + 3 = 0
(s-3)(s-1) = 0
s = 1 or s = 3
-----
If s = 1, area = 1 ; perimeter = 4
---
If s = 3, area = 9 ; perimeter = 12
--------------------
Cheers,
Stan H.

Answer by Theo(13342)

perimeter of a square = 4 * the side
area of a square = the side squared.

you get:

p = 4s
a = s^2

if the perimeter is 3 more than the area, then you get:

p = a + 3 which becomes p = s^2 + 3

you have p = s^2 + 3 and you have p = 4s

substitute s^2 + 3 for p in the second equation and you get:

s^2 + 3 = 4s

subtract 4s from both sides of this equation to get:

s^2 - 4s + 3 = 0

factor this equation to get:

(s-3) * (s+1) = 0

solve for s to get s = 3 or s = -1

s can't be negative, so s = 3

s^2 = 9

4s = 12

12 is 3 more than 9.

your solution is that the length of a side is equal to 3 and the area is equal to 9 and the perimeter is equal to 12.